An extra solar planet is found orbiting a star in the Orion nebula. Determine the mass of the star if the planet has an orbital period of 420.0 Earth Days and has an orbital radius of 1.29x1011m.


1
Expert's answer
2020-07-27T09:35:04-0400

According to the Kepler's Third Law:


"\dfrac{a^3}{T^2} = \dfrac{GM}{4\pi^2},"

the mass of the star will be:


"M = \dfrac{4a^3\pi^2}{GT^2},"

where "a = 1.29\times10^{11} m" is the orbital radius, "T = 420 days = 420\cdot 86400s = 36288000s" is the orbital period and "G = 6.6\times10^{-11} m^3kg^{-1}s^2".

Thus, obtain:


"M = \dfrac{4(1.29\times10^{11})^3\pi^2}{6.6\times10^{-11} (36288000)^2} = 9.75\times10^{29}kg"

Answer. 9.75*10^29 kg.